OpenStudy (anonymous):

Determine the midpoint given the following points A(4,2) B (-8, 18)

OpenStudy (anonymous):

(-2,10)

OpenStudy (anonymous):

it is the average of the coordinates

OpenStudy (mathteacher1729):

To verify this answer: Take a piece of graph paper and draw the two points. Draw a line between them, and find point halfway inbetween.

OpenStudy (anonymous):

-2,8

OpenStudy (anonymous):

Agreed with Satellite answer (-2,10)

OpenStudy (anonymous):

\[\frac{4+(-8)}{2}=\frac{-4}{2}=-2\] \[\frac{2+18}{2}=\frac{20}{2}=10\]

OpenStudy (anonymous):

thanks , im also giving you guys medals for helping

OpenStudy (anonymous):

ye. my bad. Satellite73 is right

OpenStudy (mathteacher1729):

See the attached picture. Of course, for midpoints of very large coordinates, it's hard to draw it all, so you'll have to use the average of the coordinates. Hopefully this will help you develop an idea of what it is we're doing with the formula.